OpenStudy (anonymous):

again, telescopic series :

OpenStudy (anonymous):

\[\frac{ 1 }{ 1010 . 2016} + \frac{ 1 }{ 1012 . 2014} + \frac{ 1 }{ 1014 . 2012} + ... + \frac{ 1 }{ 2016 . 1010} = ... ? \]

myininaya (myininaya):

the dot there means times right?

OpenStudy (anonymous):

yeah, that times...

OpenStudy (anonymous):

for this question really no idea to me

myininaya (myininaya):

well I see we can write the 1010, 1012,1014, ... as 1010+2n

myininaya (myininaya):

where that n=0,1,2,...,506

myininaya (myininaya):

to figure out the 506 i just solved 1010+2n=2016

OpenStudy (ikram002p):

http://prntscr.com/4ssja9

OpenStudy (anonymous):

can we start n = 1 ? i usually use that :)

myininaya (myininaya):

now the 2016, 2014, 2012,...,1010 as 2016-2n

myininaya (myininaya):

\[\sum_{i=0}^{506}\frac{1}{(2n+1010)(2016-2n)}\] then use partial fractions

OpenStudy (anonymous):

yeah ikram, i saw the the series has 252 pairs so just double them

myininaya (myininaya):

if you want to yes you can write it from i=1 to i=507 just write 2(n-1)+1010 and write 2016-2(n-1)

ganeshie8 (ganeshie8):

nice :) the top bound should be 503 i think...

myininaya (myininaya):

1010+2n=2016 2n=2016-1010 2n=1006 n=503 oh darn

myininaya (myininaya):

i can't solve algebraic equation

ganeshie8 (ganeshie8):

lol doesn't matter the final answer is some weird tiny number

OpenStudy (anonymous):

ha ? i i use arithmetic formula i got : 1010,1012, 1014, ..., 2016 2016 = 1010 + (n-1)2 2016 - 1010 = (n-1)2 1006 = (n-1)2 n - 1 = 503 n = 504 is tht wrong ??

myininaya (myininaya):

well you are using the formula for n=1 to n=504 we did (or i did ) n=0 to n=503

myininaya (myininaya):

yours is correct too if you choose to use 1/(1010+(n-1)*2)(2016-2(n-1)) from n=1 to n=504

OpenStudy (anonymous):

oh ok, that same :)

OpenStudy (anonymous):

well, if i start from n = 0 1/(2n + 1010)(2016-2n) = 1/2(1/2) 1/(n+505)(1008-n) = 1/4 * 1/(n+505)(1008-n) partial fractions : 1/(n+505)(1008-n) = A/(n+505) + B/(1008-n) i got A = 1/1513 and B = A = 1/1513 now i have the series : 1/4 * (1/1513)* (1/(n+505) + 1/(1008-n)) = 1/6052 * (1/(n+505) + 1/(1008-n)) what's next, @myininaya , @ganeshie8 ?

ganeshie8 (ganeshie8):

yeah im not getting ideas, i have posted it in MSE.. monitor below thread : http://math.stackexchange.com/questions/957406/dfrac-1-1010-times-2016-dfrac-1-1012-times-2014-dfrac-1

OpenStudy (ikram002p):

can we integrate ?

OpenStudy (ikram002p):

nvm lol

OpenStudy (ikram002p):

program it xD

myininaya (myininaya):

\[\frac{1}{4(1513)}[(\frac{1}{505}+\frac{1}{1008})+(\frac{1}{506}+\frac{1}{1007})+(\frac{1}{507}+\frac{1}{1006})+\cdots \\+(\frac{1}{1006}+\frac{1}{507}) +(\frac{1}{1007}+\frac{1}{506}) +(\frac{1}{1008}+\frac{1}{505}) \\ = \frac{1}{4(1513)}[\frac{2}{505}+\frac{2}{506}+\frac{2}{507}+\cdots +\frac{2}{2008}] \\ =\frac{1}{2(1513)}[\frac{1}{505}+\frac{1}{506}+\frac{1}{507}+ \cdots +\frac{1}{2008}]\]

myininaya (myininaya):

i was thinking it would work doing this oh partial fraction thing since it said it was a telescoping series

OpenStudy (ikram002p):

i think i got this one :O

myininaya (myininaya):

@ganeshie8 looks like some people replied to you

ganeshie8 (ganeshie8):

yeah still going through them, it seems it won't simplify nicely.. have to approximate only i guess